OpenStudy (anonymous):

Point A is located at (1, 5), and D is located at (−3, 7). Find the coordinates of the point that lies halfway between A and D.

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

A) (−3, 4) B) (−1, 6) C) (1, 6) D) (3, 4)

OpenStudy (anonymous):

I need help what am I doing wrong

OpenStudy (anonymous):

Calculations (x,y)=(x1+x22,y1+y22) (x,y)= (1+-3/2, 5-7/2) = -2/2, -2/2 -2/2 error -2/2 error

OpenStudy (anonymous):

use this to find mid point.. (\[(\frac{ x2-x1 }{ 2 } ,\frac{ y2-y1 }{ 2} )\]

OpenStudy (triciaal):

mid-point formula for points (x1,y1) and (x2,y2) mid point is (x1 + x2)/2 and (y1 +y2)/2 positive units

OpenStudy (anonymous):

sory..it inculed +ve sign .not -ve sign...i got confused...!!!

OpenStudy (anonymous):

https://www.desmos.com/calculator/tgkkr31hzh

OpenStudy (anonymous):

Ohh ok I see what I did wrong

OpenStudy (anonymous):

its ans is (-1,6)

OpenStudy (anonymous):

but yes, just use the midpoint formula, it is easier

OpenStudy (anonymous):

right :) Thanks Guys :)

OpenStudy (triciaal):

using one of the given coordinates add the number of units of get to the mid- point

OpenStudy (anonymous):

you are warmly welcomedd...